OpenStudy (anonymous):

How do you go about solving (√5x+6) -4 =2?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (smurfy14):

first add four to each side

OpenStudy (smurfy14):

sqrt5x+6=6 then square each side to get rid of sqrt 5x+6=36 then subtract 6 5x=30 then divide by 5 and you get x=6

OpenStudy (anonymous):

sqrt5x+6=6 sqrt5x=0 5x=0 x=0

OpenStudy (anonymous):

x = 0 is the solution

OpenStudy (anonymous):

Oh that was more simple than I was making it.

OpenStudy (anonymous):

I think I confused everyone with the way I typed it. It's more like √(5X+6) -4 =2 The root goes around 5X+6

OpenStudy (smurfy14):

\[\sqrt{5x+6}-4=2\]

OpenStudy (smurfy14):

like this? if so then my answer is correct lol

OpenStudy (anonymous):

5x+6 = 6^2 5x = 30 x= 6

OpenStudy (anonymous):

Yes that's it! I typed it incorrectly but thanks for catching that!

OpenStudy (smurfy14):

lol youre welcome! and if you dont mind could you click the blue good answer button to the far right of my name? thanks!